Steam flow Steadily Into condenser Enthalpy at entry bud exit are 2050 kj/kg and 200 kj/kg respectively. Velocity at entry and exit are 500 m/s and 10 m/s. Exit is in line with Entry . Determine Heat Transfer to cooling per unit mass of steam condensed.
1
Expert's answer
2020-10-21T06:30:46-0400

The steady flow energy equation per unit mass is:

"Q-W=H_1-H_2+\frac{(V_1^2-V_2^2)}{2}+h_2-h_1"

where

Q - Heat Transfer to cooling per unit mass,

W - external work done by system per unit mass ("W=0"),

H - Enthalpy at entry bud,

V - Velocity at entry and exit,

h - height at entry and exit ("h_2-h_1=0").


"Q=2050-200+\frac{(500^2-10^2)}{2}=1850+124950" "=126800\frac{J}{kg}"